LDG Commissioning: Performance Verification - 20 April 2010 UT

Night 2:

Observer: DMiller (Tucson) JHill (LBTO)
Telescope Operator: JHill (LBTO)
Telescope Support: MMidkiff, BCurtis
Instrument Support: JMorris

Summary

Lost the first two hours to clouds and humidity.

Were open for a couple hours of dice5 data at LDG, then closed for humidity at 6:42 UT.

Preparations

Now using TCS Build AO4 with new firmware settings on the SX M2 hexapod.

Doug has acquired the stomach flu, so he is observing from home in a semi-alive state.

Details

This is the merged observer's and operator's log as John was doing both for most of the night.

First part of night was lost to clouds and fog. We spent the time with Paul and Torsten completing the closed-dome tests for TCS Build AO4. The most significant failure was an error from TCSSendWavefront complaining about uint64_t.

Had the false CCW limit problem on the LDG rotator/chain. Escaped by holding down override while readying the rotator.

03:40 Reset encoders at Az=120 EL=90

03:45 stop GCS(s) and reboot agw-control as more than half of available swap space was used by oacserver.

04:20 Humidity down to 80% and sky is mostly clear. Still some clouds to west. Open vent doors. All-sky camera has water or something on the window even though the patio and enclosure walls are dry.

Open

04:30 Open shutter doors (with side vents closed) wind 4m/s, T=1.4 degC D=-1.1 degC

04:38 pointto M5_0384 ACQUIRE gs=0 PAM- IE=-20 CA=10

04:44 pointto M5_1210 ACQUIRE gs=0 PAM- This would have made IE=+5. Restore IE to -20

04:48 pointto BS9142 TelM=ACTIVE gs=0 - Seeing is in the 0.9 - 1.05 arcsec FWHM range, although we have just opened.

Zero out the Z12-Z21 gains as last night in AIP_IRTC_DG_L_RM_sub11.cfg

05:03 start mirror ventilation

Dice5 data on StoneI

05:07 pointto StoneIagw TelM=ACQUIRE gs=0 PAM- IE=-8

05:11 pointto StoneIagw TelM=ACTIVE gs=25 PA=0

05:15 offset 0 30

05:16 takeseq 2.0 30 irtc.20100420.051620.fits irtc.20100420_00101_cube.fits (background)

05:18 offset 0 0

05:21 takeseq 2.0 30 irtc.20100420.052118.fits 102 (on-axis star)

ignore short cubes 103 104 here

05:26 dice5 10x10 2.0 30 DETXY 1-pass Seeing is degraded to 1.5 arcsec at EL=54
irtc.20100420.052638.fits 105
irtc.20100420.052810.fits 106
irtc.20100420.052942.fits 107
irtc.20100420.053112.fits 108
irtc.20100420.053243.fits 109

05:34 dice5 10x10 2.0 30 DETXY 4-pass
irtc.20100420.053512.fits 110
irtc.20100420.053643.fits 111
irtc.20100420.053813.fits 112
irtc.20100420.053944.fits 113
irtc.20100420.054115.fits 114

irtc.20100420.054247.fits 115
- irtc.20100420.054420.fits 116
irtc.20100420.054550.fits 117
irtc.20100420.054725.fits 118
irtc.20100420.054859.fits 119

irtc.20100420.055032.fits 120
irtc.20100420.055203.fits 121
irtc.20100420.055336.fits 122
irtc.20100420.055508.fits 123
irtc.20100420.055642.fits 124

05:58 Changed ThermalLoopOffset from -10 to -12 as EL pads are running 3 degC warmer than C-ring. Stop and restart ECS. Setpoint was at -4.8 with reservoir at +0.1 (HBS has been on for much of the day)

Seeing improved to 1 arcsec for this last pass.

irtc.20100420.055815.fits 125
irtc.20100420.055944.fits 126
irtc.20100420.060116.fits 127
irtc.20100420.060249.fits 128
irtc.20100420.060421.fits 129

06:05 offset 0 30

06:06 takeseq 2.0 30 irtc.20100420.060625.fits 130 (background)

06:08 dice5 1x1 2.0 30 DETXY 5-pass - seeing improved to 0.7 -0.8 arcsec
irtc.20100420.060822.fits 131
irtc.20100420.060952.fits 132
irtc.20100420.061123.fits 133
irtc.20100420.061252.fits 134
irtc.20100420.061420.fits 135

irtc.20100420.061550.fits 136
irtc.20100420.061719.fits 137
irtc.20100420.061853.fits 138
irtc.20100420.062024.fits 139
irtc.20100420.062154.fits 140

irtc.20100420.062326.fits 141
irtc.20100420.062459.fits 142
irtc.20100420.062630.fits 143
irtc.20100420.062804.fits 144
irtc.20100420.062932.fits 145

irtc.20100420.063102.fits 146
irtc.20100420.063234.fits 147
irtc.20100420.063407.fits 148
irtc.20100420.063537.fits 149
irtc.20100420.063707.fits 150

irtc.20100420.063838.fits 151
irtc.20100420.064010.fits 152
irtc.20100420.064139.fits (shutter doors were closing during cube 153)

Close

06:42 Close for 95% humidity

07:50 Sky is clear (no fog), but humidity is hanging at 100%. Park the telescope and shutdown HBS.

-- DougMiller - 19 Apr 2010
Topic revision: r3 - 25 Apr 2010, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback